# Advancing Strategic Planning and Dynamic Control of Complex Projects

### v1

@@ 3.1 Probabilistic MCS approach
The framework enables decision-making through the MCS given the steps provided in Figure 1. (1) First, the Project network data is read, and the network $\mathcal{N}(\mathbf{x},\mathbf{y})$ is established given the project activities, risk event and any other relevant data. Note, to that extent any network planning or simulation method can be integrated, that is able to reflect the necessary logical links in the network. (2) Next, the MCS is initiated by a loop of n runs. (3) Step 3 is to sample random duration’s for the project activities, risk and other external influences (like weather). To establish the optimisation, first the type of variables must be defined. If control variables are not considered, the planning variables are defined and the optimisation initiated (continues with step 6). (4) If control variables are included, the impact $\mathbf{c_{i}}$ towards objective $i$ needs is calculated given the duration capacity of the control measure. (5) Next, the network is compiled and the delay with respect to the target duration evaluated. In case of a delay, the optimisation is initiated. (6) The optimisation is run by defining the objective functions, the preference functions and their related weights, in accordance with the prior established mathematical formulation for the optimisation (see section 2). Once defined, the optimisation is performed utilising the IMAP method. To arrive at the highest aggregated preference score $A$, the values of the objectives functions are calculated and with that the corresponding value of the preference functions. Once established, the aggregated preference score can be determined according to the principles of references aggregation with preference maximisation (more information see the section on the optimisation below). (7) After every iteration of the MCS the result is stored. (8) Once the MCS reaches the $n^{th}$ iteration, the results of the individual optimisations can be presented with the criticality index (how often does a certain variable or a combination of variables occur). Given the collected data, detailed statistical analysis are enabled.
@@ 3.2 Integrative multi-objective optimisation approach
To integrate the human goal-oriented behaviour in project management, and associative decision-making the multi-objective optimisation formulation (defined in section 2) is integrated into the MCS as part of the IMAP optimisation method. The optimisation is executed within every iteration of the probabilistic MCS, aiming to select the most effective set of project planning variables and or control measures given the specific realisation of the network within that iteration.
@@ 4 Demonstrative planning & control applications
4.1 Strategic planning application
@@ 4.1.1 Performance functions
where the physical variables $\mathbf{y}$ contain the information on the respective vessel deck space. To ensure that the project is executed by at least one vessel, the sum of all vessels on the project must be greater than one. The technical feasibility is guaranteed by a constraint ensuring the anchor’s resistance meets or exceeds the applied forces defined by further performance functions ($F_{4}$ ,…, $F_{7}$). For a thorough exposition of these formulations and their underlying principles, the reader is referred to Wolfert (2023) or Van Heukelum et al. (2024).
@@ 4.1.2 Objective functions
The optimisation framework considers the following four objectives that form the link between the network performance function and the preference functions. For a thorough exposition of these formulations and their underlying principles, the reader is referred to Wolfert (2023) or Van Heukelum et al. (2024). The following objective functions are rewritten based upon to suit the needs of the Odycon framework.
@@ 4.1.4 Results
Stakeholder $k$ $w^{\prime}_{k,PD}$ $w^{\prime}_{k,S}$ $w^{\prime}_{k,C}$ $w^{\prime}_{k,PD}$ $w_{k}$
@@
Figure 3: Criticality index of number of small OCV’s (top left), large OCV’s (top right), barges (bottom left) and vessel combinations (bottom right) for both optimisation cases. Note, only combinations that stored during the simulation & optimization are considered in this figure.
4.2 Dynamic control application
@@ 6 Conclusion
First, the complex reality of project management is addressed by identification of the most effective and desired set of project planning & control variables given the uncertainties and stochastic factors while considering overall group satisfaction. The interplay with stakeholder-oriented concurrent objectives is addressed by the integration of the ’associative preference domain’, enabling an optimal solution that best fits the common purpose, using a-priori optimization. This approach supports satisfactory project planning and control, allowing stakeholders to meet their individual needs through collective performance. The model results of both demonstrative applications show that joint project success, in alignment with the defined common purpose, will increase when individual stakeholders relinquish their pure self-interest, ultimately enabling an optimal solution that benefits the whole group. Here, Odycon aligns with previous empirical results of management studies, as mentioned in the introduction. The developed methodology allows for the integration and combination of variables from different project phases, as well as flexible adaptation and development tailored to project and stakeholder specific needs.

### v2

@@ 3.1 Probabilistic MCS approach
The framework enables decision-making through the MCS given the steps provided in Figure 1. (1) First, the Project network data is read, and the network $\mathcal{N}(\mathbf{x},\mathbf{y})$ is established given the project activities, risk event and any other relevant data. Note, to that extent any network planning or simulation method can be integrated, that is able to reflect the necessary logical links in the network. (2) Next, the MCS is initiated by a loop of n runs. (3) Step 3 is to sample random duration’s for the project activities, risk and other external influences (like weather). To establish the optimisation, first the type of variables must be defined. If control variables are not considered, the planning variables are defined and the optimisation initiated (continues with step 6). (4) If control variables are included, the impact $\mathbf{c_{i}}$ towards objective $i$ needs is calculated given the duration capacity of the control measure. (5) Next, the network is compiled and the delay with respect to the target duration evaluated. In case of a delay, the optimisation is initiated. (6) The optimisation is run by defining the objective functions, the preference functions (which defines the desired outcomes for the objective functions) and their related weights, in accordance with the prior established mathematical formulation for the optimisation (see section 2). Once defined, the optimisation is performed utilising the IMAP method. To arrive at the highest aggregated preference score $A$, the values of the objectives functions are calculated and with that the corresponding value of the preference functions. Once established, the aggregated preference score can be determined according to the principles of references aggregation with preference maximisation (more information see the section on the optimisation below). (7) After every iteration of the MCS the result is stored. (8) Once the MCS reaches the $n^{th}$ iteration, the results of the individual optimisations can be presented with the criticality index (how often does a certain variable or a combination of variables occur). Given the collected data, detailed statistical analysis are enabled.
@@ 3.2 Integrative multi-objective optimisation approach
To integrate the human goal-oriented behaviour in project management, and associative decision-making the multi-objective optimisation formulation (defined in section 2) is integrated into the MCS as part of the IMAP optimisation method. The optimisation is executed within every iteration of the probabilistic MCS, aiming to select the most effective set of project planning variables and or control measures given the specific realisation of the network within that iteration. The
@@ 4 Demonstrative planning & control applications
4.1 Strategic Planning application : Agree First & Act Feasibly
@@ 4.1.1 Performance functions
where the variables $\mathbf{y}$ contain the information on the respective vessel deck space or are related to the risks (and duration of the activity). To ensure that the project is executed by at least one vessel, the sum of all vessels on the project must be greater than one. The technical feasibility is guaranteed by a constraint ensuring the anchor’s resistance meets or exceeds the applied forces defined by further performance functions ($F_{4}$ ,…, $F_{7}$). For a complete exposition of these formulations and their underlying principles, the reader is referred to Van Heukelum et al. (2024).
@@ 4.1.2 Objective functions
The optimisation framework considers the following four objectives that form the link between the network performance function and the preference functions. For a complete exposition of these formulations and their underlying principles, the reader is referred to Wolfert (2023) or Van Heukelum et al. (2024). The following objective functions are rewritten based upon to suit the needs of the Odycon framework.
@@ 4.1.4 Results
Stakeholder $k$ $w^{\prime}_{k,PD}$ $w^{\prime}_{k,S}$ $w^{\prime}_{k,C}$ $w^{\prime}_{k,F}$ $w_{k}$
@@
Figure 3: Criticality index of number of small OCV’s (top left), large OCV’s (top right), barges (bottom left) and vessel combinations (bottom right) for both optimisation cases. Note, only combinations that stored during the simulation & optimisation are considered in this figure.
4.2 Dynamic Control application: Adapt Flexibly
@@ 6 Conclusion
First, the complex reality of project management is addressed by identification of the most effective and desired set of project planning & control variables given the uncertainties and stochastic factors while considering overall group satisfaction. The interplay with stakeholder-oriented concurrent objectives is addressed by the integration of the ’associative preference domain’, enabling an optimal solution that best fits the common purpose, using a-priori optimisation. This approach supports satisfactory project planning and control, allowing stakeholders to meet their individual needs through collective performance. The model results of both demonstrative applications show that joint project success, in alignment with the defined common purpose, will increase when individual stakeholders relinquish their pure self-interest, ultimately enabling an optimal solution that benefits the whole group. Here, Odycon aligns with previous empirical results of management studies, as mentioned in the introduction. The developed methodology allows for the integration and combination of variables from different project phases, as well as flexible adaptation and development tailored to project and stakeholder specific needs.
@@
Acknowledgements
The authors would like to thank Dr. Ruud Binnekamp of Delft University of Technology for his valuable reflections and guidance in completing this paper.